find x and y

x + 2y = 10
3x - y = 2

Answer by checkley75(3666)

x+2y=10
x=10-2y now substitute (10-2y) for x in the other equation & solve for y
3(10-2y)-y=2
30-6y-y=2
-7y=-28
y=-28/-7
y=4 answer.
x+2*4=10
x+8=10
x=10-8
x=2 answer.